OpenStudy (anonymous):

what are the factors if the expression x^2-10x-24 ? please solve

OpenStudy (anonymous):

(x+2)(x-12)

OpenStudy (anonymous):

do you know how to solve

OpenStudy (anonymous):

Do you need the solving for finding them?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Give me one second.

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

For a polynomial of the form ax^2+bx+c, pull out the values of a,b, and c. Solve by grouping, using two factors of a*c(-24) whose sum is b(-10). \[x^2-10x-24\] \[a = 1, b = -10, c = -24\] Contiuing, the middle term as a sum of two terms whose product is: 1*-24=-24 and whose sum is =10. \[(x^2-12x+2x-24)\] Factor the GCF (greatest common factor) from each group. \[(x(x-12)=2(x-12))\] Factor the polynomial by grouping the first two terms together and finding the GCF. Next, group the second two terms together and find the GCF. Since bother groups contain the factor (x-12), they can be combined. Resulting in: \[(x+2)(x-12)\]